18 . Jase Manufacturing Co.'s static budget for 7,900 units of production includes $39,500 for direct labor and $2,370 for variable electric power. Total fixed costs are $37,700. At 10,400 units of production, a flexible budget would show

a.variable costs of $55,120 and $49,630 of fixed costs

b.variable and fixed costs totaling $79,570

c.variable costs of $55,120 and $37,700 of fixed costs

d.variable costs of $41,870 and $37,700 of fixed costs
Answer #1

Direct labor per unit = $39,500 / 7,900 = $5

Variable electric power per unit = $2,370 / 7,900 = $0.3

At 10,400 units, Variable costs = 10,400 * ($5 + $0.3) = $55,120

Fixed costs = $37,700

The answer is C.